If you have a parallelogram with two lines bisecting. with m+9.1 on one half of the bisector and 4k-3.5 on the other side. and on the other line there is 2.6k and m. How do you solve this without having to deal with 2 variables?
Answer by richwmiller(17219)

Let's look at what we know before we panic.
Since the line bisects it creates two equal parts.
m+9.1=4k-3.5
oh no we have one equation with two variables.
Stay calm; we aren't finished yet.
2.6k=m
Oh now we have two equations with two variables. So we can solve the two equations and find out where they intersect.
Well, look at that m=2.6k
We can substitute 2.6k for m in the first equation.
2.6k+9.1=4k-3.5
We now have one equation with one variable.
Panic attack avoided.
k=9
2.6k=m
2.6*(9)=m
23.4 = m
We can check.
23.4+9.1=4(9)-3.5
32.5=32.5
ok